Hello! Can someone explain how to solve this problem? I could not solve it and need to in order to move forward with the assignment.
Thanks!

Texas-Q Company produces and sells barbeque grills. Texas-Q sells three models: a small portable gas grill, a larger stationaRefer to the list below for the exact wording of text items within your income statement. Amount Descriptions Operating incom1. What is the sales mix of portable grills to stationary grills to smokers? 2. Compute the break-even quantity of each produ3(b) Prepare an income statement for Texas-Q for the coming year. Refer to the list of Amount Descriptions for the exact word4. Compute the margin of safety for the coming year. The margin of safety for the coming year is $
Answer #1

1.

Portable Grills Stationery Grills Smokers
Sell 20800 52000 5200
Sales Mix 4 10 1
Sales Mix of Portable grills to stationery grills to smokers is 4:10:1

2.

Portable Grills 8720
Stationery Grills 21800
Smokers 2180

Working:

Particular Portable Grills Stationery Grills Smokers Total
Selling Price per unit $                    86.00 $                  200.00 $                252.00
Variable Cost per unit $                    44.00 $                  128.00 $                139.00
Contribution per unit $                    42.00 $                    72.00 $                113.00
Sales Mix 4 10 1
Total Contribution Margin $                 168.00 $                  720.00 $                113.00 $            1,001.00
Break even unit = Fixed Cost / Contribution Margin
Break even unit = 2182180 /1001
Break even units = 2180 units
Particular Portable Grills Stationery Grills Smokers
Break Even units 2180 2180 2180
Sales Mix 4 10 1
Break even quantity 8720 21800 2180

3a.

Contribution Margin Ratio = 38.56%
Break Even Revenue = $5,659,180

Working:

Particular   Portable Grills   Stationery Grills   Smokers   Total
Sales Revenue $      1,788,800.00 $     10,400,000.00 $     1,310,400.00 $   13,499,200.00
Less: Variable Cost   $          915,200.00 $       6,656,000.00 $        722,800.00 $     8,294,000.00
Contribution Margin   $          873,600.00 $       3,744,000.00 $        587,600.00 $     5,205,200.00
Less: Fixed Cost   $     2,182,180.00
Operating Income   $    3,023,020.00
Contribution Margin Ratio = Contribution Margin / Sales Revenue  
Contribution Margin Ratio =$5205,200 / 13,499,200 *100
Contribution Margin Ratio = 38.56%
Break Even Revenue = Fixed Cost / Contribution Margin Ratio
Break Even Revenue = 2182180/38.56%
Break Even Revenue = $5,659,180

3b.

Income Statement
Sales Revenue $    13,499,200.00
Less: Variable Cost   $      8,294,000.00
Contribution Margin   $      5,205,200.00
Less: Fixed Cost   $      2,182,180.00
Operating Income   $      3,023,020.00

4.

Margin of Safety = Sales - Break even Sale Revenue
Margin of Safety = 13,499,200 - 5,659,180
Margin of Safety = $7840020